# Robust Point Matching with Distance Profiles

### v2

@@ 4 Simulations
We first provide a simulation study to demonstrate the performance of the matching procedure discussed in Section 2. Figure 2 shows the matching accuracy defined as follows:
@@
where $t,s$ are defined as in Theorem 1. In other words, we compute the proportion of correctly matched pairs among $X_{i}$’s from the first $K$ components. This accuracy measure is less strict and thus can be more practically relevant than the matching guarantee in Theorem 1 that is the probability of matching without any error. For each plot of Figure 2, we compare the case where there is one outlier, that is, $t=s=K+1$ in Assumption 1, and the baseline case where there is no outlier, namely, $t=s=K$. In the former case, $\Omega$ defined in (2.2) is $0.0345$, while $\Omega=\min_{\alpha\neq\beta}\overline{W}(\alpha,\beta)=0.1144$ for the later case as we have $\overline{W}(\alpha,\alpha)=0$ for all $\alpha\in[K]$. For both cases, we let $\sigma=\sigma_{k}=\tau_{k}$ for all $k\in[t]$ in Assumption 1 and vary $\sigma$. We can observe that the matching accuracy deteriorates as $\sigma$ grows for both cases. Though the presence of an outlier deteriorates the matching accuracy as expected, we can see that the proposed method is robust to the outlier in that the accuracy is not significantly lower than the baseline case considering the signal-to-noise ratio ($\frac{1}{K}=0.1$).
(a) $n=200$
(b) $n=500$
(c) $n=1000$
Figure 2: Accuracy of Gaussian mixture matching with $d=3$, $K=10$, and the sample size $n=m\in\{200,500,1000\}$. For each plot, the blue solid line amounts to the case where there is one outlier, that is, $t=s=K+1$ in Assumption 1, while the red dashed line is the baseline case where there is no outlier, namely, $t=s=K$. We let $\sigma_{k}=\tau_{k}=\sigma$ for all $k\in[t]$ and vary $\sigma$. For each $\sigma$, the sampling from $\mu,\nu$ is repeated 100 times and the marked points show the average of the matching accuracy over 100 realizations, while the shaded regions show the standard deviations.
@@ 5 Conclusion
In this paper, we have proposed a matching procedure based on distance profiles and established its robustness. The proposed procedure is easily implementable without having to solve nontrivial optimization problems over permutations. Most importantly, we have shown that the proposed method is provably robust to outliers and noise under some model assumptions that we believe capture essential aspects of empirical applications. To the best of our knowledge, this is the first work that provides theoretical guarantees for the robustness of a matching procedure based on distance profiles in the context of object matching. We have also demonstrated the empirical performance of the proposed method through real data and simulations.
We suggest several directions for future research. On the theoretical side, it would be interesting to consider other types of matching guarantees in Theorem 1. As mentioned in Section 4, controlling the matching accuracy (4.1) may be practically more useful. For any small $\eta\in(0,1)$, showing that the matching accuracy (4.1) is larger than $1-\eta$ with high probability would be a meaningful result. Similarly, for Theorem 2, it would be interesting to consider the probability that $\hat{\pi}$ is close to $\pi^{\ast}$ under a certain discrepancy, such as the Hamming distance, instead of the probability of a perfect recovery $\hat{\pi}=\pi^{\ast}$. On the practical side, it would be interesting to delve deeper into the procedure of finding inliers, that is, the points from the common part of two objects. In Remark 1, though we have noted that the threshold $\rho$ can be chosen to identify the inliers at least in theory, developing a practical method applicable to real data is very challenging in general. We leave these directions for future research.
@@ A.1 Details of Remark 1
The proof of (2.6) in Remark 1 is similar to the proof of Theorem 1. First, recall that the event $C=(2R(Z_{1}+Z_{2})+2(U+V)<\Omega)$ implies $s(\pi(i))=t(i)$ for any $i\in[n]$ such that $t(i)\in[K]$. Hence, on the event $C$, one can deduce that
$\max_{i\in[n]:t(i)\in[K]}W(i,\pi(i))\leq\max_{\alpha\in[K]}\overline{W}(\alpha,\alpha)+R(Z_{1}+Z_{2})+(U+V).$
@@
$\begin{split}\min_{i\in[n]:t(i)\notin[K]}W(i,\pi(i))&=\min_{i\in[n]:t(i)\notin[K]}\min_{j\in[m]}W(i,j)\\ &\geq\min_{\alpha\in[t]\backslash[K],\beta\in[s]}\overline{W}(\alpha,\beta)-R(Z_{1}+Z_{2})-(U+V).\end{split}$
Now, let $C^{\prime}:=(2R(Z_{1}+Z_{2})+2(U+V)<\Omega^{\prime})$. From (2.2) and (2.5), we have $\Omega^{\prime}\leq\Omega$ and thus $C^{\prime}\subset C$. Therefore, on the event $C^{\prime}$, we have
$\min_{i\in[n]:t(i)\notin[K]}W(i,\pi(i))-\max_{i\in[n]:t(i)\in[K]}W(i,\pi(i))\geq\Omega^{\prime}-2R(Z_{1}+Z_{2})-2(U+V)>0.$
Now, we have $\mathbb{P}(C^{\prime})\geq 1-\delta$ by repeating the proof of Theorem 1 with $\Omega$ in (2.3), (2.4) replaced by $\Omega^{\prime}$. Therefore, we have (2.6) with probability at least $1-\delta$.
@@ A.2 Lemmas
We have used the following lemmas on the concentration of the multinomial distribution and the chi-squared distribution in the proofs of Theorems 1 and 2. The former is from [Devroye, 1983], while the latter is a standard result that can be deduced from the Chernoff bound; for instance, see Chapter 2 of [Vershynin, 2018].
Lemma 1 (Lemma 3 of [Devroye, 1983]).
Suppose $(z_{1},\ldots,z_{t})$ is from the multinomial distribution with the number of trials $n$ and the probability vector $(p_{1},\ldots,p_{t})\in\Delta_{t}$. Then, for any $\delta\in(0,1)$ such that $\delta\leq 3e^{-4t/5}$, we have
$\mathbb{P}\left(\sum_{k=1}^{t}\left|\frac{z_{k}}{n}-p_{k}\right|\geq 5\sqrt{\frac{\log(3/\delta)}{n}}\right)\leq\delta.$

### v5

@@ 5 Simulations
5.1 Gaussian Mixture Matching
We first provide a simulation study to demonstrate the performance of the matching procedure discussed in Section 2. Figure 3(a) shows the probability of the perfect matching, namely, the left-hand side of (2.2). Here, $d=3$ and $K=10$, and we consider the clean case $t=s=K$ and the case with one outlier $t=s=K+1$ for the sample size $n=m\in\{500,1000\}$; also, we let $p_{k}=q_{k}=1/t$ and $\sigma=\sigma_{k}=\tau_{k}$ for all $k\in[t]$, and vary $\sigma$. We can observe that the probability of the perfect matching decreases as $\sigma$ grows for both cases. Though the presence of an outlier deteriorates the matching performance as expected, we can verify the outlier robustness of the proposed method in that the recovery probability is sufficiently high when $\sigma$ is small. Comparing the plots for $n=m=500$ and $n=m=1000$, we can see that the matching performance is better for the larger sample size as it better approximates the population-level distance profiles.
Although the analysis in Section 2 was based on the probability of the perfect matching, it is reasonable to expect matching errors in practice. Figure 3(b) shows the matching accuracy for the same setting as Figure 3(a) with $n=m=1000$, where the matching accuracy is the proportion of correctly matched $X_{i}$’s from the first $K$ components:
@@ 5.1 Gaussian Mixture Matching
where $t,s$ are defined as in Theorem 1. Again, we can see that the proposed method performs reasonably well. Figure 3(c) compares the case with one outlier ($K=10$ and $t=s=11$) to the case with more outliers but with the same outlier proportion ($K=20$ and $t=s=22$) and to the case with more outliers with larger outlier proportions ($K=20$ and $t=s=24$). We can observe that the matching accuracy decreases as the outlier proportion increases, which is consistent with the theoretical results in Section 2. Even when the outlier proportion $(t-K)/K$ is the same, the matching accuracy is lower for larger $t$ as the larger effective dimensions of the underlying distributions $\mu,\nu$ lead to larger estimation errors given the same sample size. Figure 3(d) compares the case with $d=3$ to the case with $d=10$ for $K=10$ and $t=s=11$. We can see that the matching accuracy is lower and deteriorates much faster for larger $d$.
(a) Perfect matching probability
(b) Matching accuracy ($K=10$)
(c) Matching accuracy (varying $K$)
(d) Matching accuracy (varying $d$)
Figure 3: Gaussian mixture matching results. (a) shows the probability of the perfect matching shown on the left-hand side of (2.2) with $d=3$, $K=10$, and the sample size $n=m\in\{500,1000\}$; we let $p_{k}=q_{k}=1/t$ and $\sigma=\sigma_{k}=\tau_{k}$ for all $k\in[t]$, and vary $\sigma$. (b) shows the matching accuracy defined in (5.1) for the same setting as (a) with $n=m=1000$. (c) and (d) compare the matching accuracy for different settings with varying $K$ or $d$, while we keep the same sample size $n=m=1000$. For all plots, the sampling from $\mu,\nu$ is repeated 100 times. For (a), the marked points show the probability estimated from the 100 realizations. For (b)-(d), the marked points show the average of the matching accuracy over 100 realizations, while the shaded regions show the standard deviations. The blue solid lines of (b)-(d) are the same, which correspond to the case with $K=10$ and $t=s=11$.
5.2 Noise Stability
@@ 6 Conclusion
We have proposed a matching procedure based on distance profiles, which is easily implementable without having to solve nontrivial optimization problems over permutations. Most importantly, we have shown that the proposed method is provably robust to outliers and noise under some model assumptions that we believe capture essential aspects of empirical applications. To the best of our knowledge, this is the first work that provides theoretical guarantees for the robustness of a matching procedure based on distance profiles in the context of object matching. Also, we have extended the concept of distance profiles to the abstract setting and connected the proposed matching procedure to the Gromov-Wasserstein distance and its lower bound, with a new sample complexity result derived based on the properties of distance profiles. We have also demonstrated the empirical performance of the proposed method through real data and simulations.
We suggest several directions for future research. On the theoretical side, it would be interesting to consider other types of matching guarantees in Theorem 1. As mentioned in Section 5, the matching accuracy (5.1) may be practically more useful. Similarly, for Theorem 2, it would be interesting to consider the probability that $\hat{\pi}$ is close to $\pi^{\ast}$ under a certain discrepancy, such as the Hamming distance, instead of the probability of $\hat{\pi}=\pi^{\ast}$. On the practical side, it would be interesting to delve deeper into the procedure of finding inliers, that is, the points from the common part of two objects. In Remark 2, though we have noted that the threshold $\rho$ can be chosen to identify the inliers at least in theory, developing a practical method applicable to real data is very challenging in general. Lastly, building on the sample complexity result in Section 4, we may analyze the outlier robustness of the third lower bound in general cases. We leave these directions for future research.
Acknowledgments
The work of YK was partially supported by DMS-2339439, DE-SC0022232, and Sloan research fellowship. The work of YH was partially supported by DE-SC0022232.
@@ A.1 Details of Remark 2
First, recall from the proof of Theorem 1 that the event $C=(2R(Z_{1}+Z_{2})+2(U+V)<\Omega)$ implies $s(\pi(i))=t(i)$ for any $i\in[n]$ such that $t(i)\in[K]$. Hence, on the event $C$, one can deduce that for such $i\in[n]$, we have
$W(i,\pi(i))\leq\overline{W}(t(i),t(i))+R(Z_{1}+Z_{2})+U+V,$
which follows from $|W(i,j)-\overline{W}(t(i),s(j))|\leq R(Z_{1}+Z_{2})+(U+V)$ for any $i,j\in[n]$. Therefore, invoking Lemma 1, we have
$\max_{\begin{subarray}{c}i\in[n]\\ t(i)\in[K]\end{subarray}}W(i,\pi(i))\leq R\cdot\bigg{(}1-\sum_{k\in[K]}p_{k}\bigg{)}+R(Z_{1}+Z_{2})+(U+V).$
@@
$\begin{split}\min_{\begin{subarray}{c}i\in[n]\\ t(i)\notin[K]\end{subarray}}W(i,\pi(i))&\geq\min_{\begin{subarray}{c}i\in[n]\\ t(i)\notin[K]\end{subarray}}\overline{W}(t(i),s(\pi(i)))-R(Z_{1}+Z_{2})-(U+V)\\ &\geq\min_{\begin{subarray}{c}\alpha\in[t]\backslash[K]\\ \beta\in[s]\end{subarray}}\overline{W}(\alpha,\beta)-R(Z_{1}+Z_{2})-(U+V).\end{split}$
Now, let $C^{\prime}$ be the event $(2R(Z_{1}+Z_{2})+2(U+V)<\Omega^{\prime})$, where $\Omega^{\prime}$ represents the quantity in (2.10). Clearly, we have $\Omega^{\prime}\leq\Omega$ and thus $C^{\prime}\subset C$. Therefore, on the event $C^{\prime}$, we have
$\min_{\begin{subarray}{c}i\in[n]\\ t(i)\notin[K]\end{subarray}}W(i,\pi(i))-\max_{\begin{subarray}{c}i\in[n]\\ t(i)\in[K]\end{subarray}}W(i,\pi(i))\geq\Omega^{\prime}-2R(Z_{1}+Z_{2})-2(U+V)>0.$
Now, we have $\mathbb{P}(C^{\prime})\geq 1-\delta$ by repeating the proof of Theorem 1 with the left-hand side of (2.3), which is $\Omega$, replaced by $\Omega^{\prime}$. Therefore, we have (2.11) with probability at least $1-\delta$.
@@ A.2 Lemmas
We have used the following lemmas on the concentration of the multinomial distribution and the chi-squared distribution in the proofs of Theorems 1 and 2. The former is an application of Proposition A.6.6 of [36], while the latter is a standard result that can be deduced from the Chernoff bound; for instance, see Chapter 2 of [38].
Lemma 3.
Suppose $(z_{1},\ldots,z_{t})$ is from the multinomial distribution with the number of trials $n$ and the probability vector $(p_{1},\ldots,p_{t})\in\Delta_{t}$. Let $\mathcal{A}\subset 2^{[t]}$ be a collection of subsets of $[t]$. Then, for any $\varepsilon\in(0,1)$, we have
$\mathbb{P}\left(\max_{A\in\mathcal{A}}\left|\sum_{k\in A}\left(\frac{z_{k}}{n}-p_{k}\right)\right|\geq\varepsilon\right)\leq 2|\mathcal{A}|\cdot e^{-2n\varepsilon^{2}}.$
@@
A.3 Extension of Theorem 4 to General Metric Measure Spaces
From (4.11) in the proof of Theorem 4, one can deduce that Theorem 4 can be extended to any metric measure space $(\mathcal{X},d_{\mathcal{X}},\mu)$ by replacing $d+1$ in (4.10) with the VC dimension of $\mathcal{F}_{\mathcal{X}}$ defined in the proof, which is the VC dimension of the set of all balls in $\mathcal{X}$. Unfortunately, controlling the VC dimension of the set of all balls in a general metric measure space is highly nontrivial. A more straightforward quantity is the $\varepsilon$-covering number $\mathcal{N}(\varepsilon,\mathcal{X},d_{\mathcal{X}})$, namely, the smallest possible number of $\varepsilon$-balls to cover $\mathcal{X}$. The following theorem derives an upper bound using the covering number for a general bounded metric measure space $(\mathcal{X},d_{\mathcal{X}},\mu)$.
Theorem 5.
Let $(\mathcal{X},d_{\mathcal{X}},\mu)$ be a metric measure space and $\hat{\mu}=\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}}$ be the empirical measure based on $X_{1},\ldots,X_{n}$ that are i.i.d. from $\mu$. Suppose $\Delta=\sup_{x,x^{\prime}\in\mathcal{X}}d_{\mathcal{X}}(x,x^{\prime})<\infty$ and the support of $\mu$ is exactly $\mathcal{X}$. Let $\mathcal{N}(\varepsilon,\mathcal{X},d_{\mathcal{X}})$ be the $\varepsilon$-covering number of $\mathcal{X}$ under $d_{\mathcal{X}}$. Then,
$\mathbb{E}\left[\sup_{x\in\mathcal{X}}W_{1}(\mu_{x},\hat{\mu}_{x})\right]\leq\frac{C}{\sqrt{n}}\int_{0}^{\Delta/4}\sqrt{\frac{\Delta}{\varepsilon}+\log\mathcal{N}(\varepsilon,\mathcal{X},d_{\mathcal{X}})}\,\mathrm{d}\varepsilon,$ (A.1)
where $C$ is an absolute constant that is independent of $n$, $\mathcal{X}$, and $\mu$.
Proof.
By the Kantorovich-Rubinstein theorem (Theorem 1.14 of [39]), we have
$W_{1}(\mu_{x},\hat{\mu}_{x})=\sup_{f\in\mathrm{Lip}_{1}}\left(\mu_{x}f-\hat{\mu}_{x}f\right),$
[…]
$\sup_{x\in\mathcal{X}}W_{1}(\mu_{x},\hat{\mu}_{x})=\sup_{x\in\mathcal{X}}\sup_{f\in\mathrm{Lip}_{1}}\left(\mu_{x}f-\hat{\mu}_{x}f\right)=\sup_{\phi\in\mathcal{F}}\left(\mu\phi-\hat{\mu}\phi\right),$
where =$\mathcal{F}=\{f\circ d_{\mathcal{X},x}:f\in\mathrm{Lip}_{1},x\in\mathcal{X}\}$ which is a function class from $\mathcal{X}$ to $\mathbb{R}$. Here, $d_{\mathcal{X},x}\colon\mathcal{X}\to\mathbb{R}$ is defined by $d_{\mathcal{X},x}(\cdot)=d_{\mathcal{X}}(x,\cdot)$. Using symmetrization and Dudley’s chaining argument (see Chapter 8 of [38]), one can derive the following:
$\mathbb{E}\left[\sup_{x\in\mathcal{X}}W_{1}(\mu_{x},\hat{\mu}_{x})\right]=\mathbb{E}\left[\sup_{\phi\in\mathcal{F}}\left(\mu\phi-\hat{\mu}\phi\right)\right]\leq\frac{C}{\sqrt{n}}\int_{0}^{\Delta/2}\sqrt{\log\mathcal{N}(\varepsilon,\mathcal{F},\|\cdot\|_{\mathcal{X}})}\,\mathrm{d}\varepsilon,$ (A.2)
where $\mathcal{N}(\varepsilon,\mathcal{F},\|\cdot\|_{\mathcal{X}})$ denotes the $\varepsilon$-covering number of $\mathcal{F}$ with respect to $\|\cdot\|_{\mathcal{X}}$, namely, the sup norm defined by $\|\phi\|_{\mathcal{X}}=\sup_{x\in\mathcal{X}}|\phi(x)|$ for any $\phi\colon\mathcal{X}\to\mathbb{R}$. Here, we are using the fact that $\sup_{\phi\in\mathcal{F}}\|\phi\|_{\mathcal{X}}\leq\Delta$ since $\sup_{z\in[0,\Delta]}|f(z)|\leq\Delta$ for any $f\in\mathrm{Lip}_{1}$ by Lipschitzness. Now, noticing that $\mathcal{F}$ is a function class consisting of composite functions, we upper bound $\mathcal{N}(\varepsilon,\mathcal{F},\|\cdot\|_{\mathcal{X}})$ by the product of the $\varepsilon$-covering number of $\mathrm{Lip}_{1}$ with respect to the sup norm $\|\cdot\|_{\infty}$ on $[0,\Delta]$, that is, $\|f\|_{\infty}=\sup_{z\in[0,\Delta]}|f(z)|$, and the $\varepsilon$-covering number of $\mathcal{X}$ with respect to $d_{\mathcal{X}}$. Concretely, we claim
$\mathcal{N}(2\varepsilon,\mathcal{F},\|\cdot\|_{\mathcal{X}})\leq\mathcal{N}(\varepsilon,\mathrm{Lip}_{1},\|\cdot\|_{\infty})\times\mathcal{N}(\varepsilon,\mathcal{X},d_{\mathcal{X}}).$ (A.3)
To see this, let $\mathrm{Lip}_{1}^{\varepsilon}\subset\mathrm{Lip}_{1}$ and $\mathcal{X}^{\varepsilon}\subset\mathcal{X}$ be $\varepsilon$-coverings of $\mathrm{Lip}_{1}$ and $\mathcal{X}$, respectively. For $f\in\mathrm{Lip}_{1}$ and $x\in\mathcal{X}$, we can find $f^{\prime}\in\mathrm{Lip}_{1}^{\varepsilon}$ and $x^{\prime}\in\mathcal{X}^{\varepsilon}$ such that $\|f-f^{\prime}\|_{\infty}\leq\varepsilon$ and $d_{\mathcal{X}}(x,x^{\prime})\leq\varepsilon$. Then,
$\begin{split}\|f\circ d_{\mathcal{X},x}-f^{\prime}\circ d_{\mathcal{X},x^{\prime}}\|_{\mathcal{X}}&\leq\|f\circ d_{\mathcal{X},x}-f\circ d_{\mathcal{X},x^{\prime}}\|_{\mathcal{X}}+\|f\circ d_{\mathcal{X},x^{\prime}}-f^{\prime}\circ d_{\mathcal{X},x^{\prime}}\|_{\mathcal{X}}\\ &\leq\|d_{\mathcal{X},x}-d_{\mathcal{X},x^{\prime}}\|_{\mathcal{X}}+\|f-f^{\prime}\|_{\infty}\\ &\leq d_{\mathcal{X}}(x,x^{\prime})+\varepsilon\\ &\leq 2\varepsilon.\end{split}$
Hence, we obtain (A.3). The standard upper bound on the covering number of $\mathrm{Lip}_{1}$ (see Exercise 8.2.7 of [38]) tells
$\log\mathcal{N}(\varepsilon,\mathrm{Lip}_{1},\|\cdot\|_{\infty})\leq\frac{c\Delta}{\varepsilon},$ (A.4)
where $c$ is some absolute constant. Combining (A.2), (A.3), and (A.4), we obtain (A.1). ∎
Remark 3.
If one assumes that $\mathcal{X}$ is a compact subset of $\mathbb{R}^{d}$ and $d_{\mathcal{X}}$ is the standard Euclidean distance in Theorem 5, the bounds coincide with those in Theorem 4. To see this, it suffices to notice that $\mathcal{N}(\varepsilon,\mathcal{X},d_{\mathcal{X}})$ scales as $(1+2/\varepsilon)^{d}$; see Section 4.2 of [38]. The key difference, however, is that Theorem 5 requires compactness of the support of $\mu$, while Theorem 4 does not.